# Hume: Introducing System-2 Thinking in Visual-Language-Action Model

### v1

@@ Abstract
Abstract
This supplementary material accompanies the main paper by providing more detailed visualization analysis of Hume’s workflow, as well as implementation details and additional experimental results:
$\triangleright$ Sec. 6: Detail Hume’s workflow by visualize the value-guided thinking and cascaded action denoising processes.
$\triangleright$ Sec. 7: Video demonstration and anonymous link: https://hume-vla.github.io.
$\triangleright$ Sec. 8: Implementation details including loss functions and hyperparameters.
$\triangleright$ Sec. 9: Experimental details in simulation and real world including experiment setup and detailed results.
@@ 3.1 Value-Guided System-2 Thinking
State-Action Value Estimation The state-action values are estimated with the proposed value-query head built on the same VLM backbone via learning a latent conditioned Q function. The value-query head is composed of two critic networks estimating state-action values and one actor network for assisting the training of critic networks. Specifically, a special query token $\mathbf{q}_{t}$ is introduced and attached at the end of the VLM input sequence, which is a learnable token with the same embedding dimension as the language tokens. Then, for one action chunk $\mathbf{A}_{t}$ (either ground-truth action $\mathbf{A}_{t}$ or denoised candidate actions $\mathbf{A}_{t}^{\tau_{n}}$), it is combined with this special query token $\mathbf{q}_{t}$ to feed into the value-query head. Due to its last position of the input sequence, the query token $\mathbf{q}_{t}$ attends to all previous tokens and aggregates necessary information from the VLM inputs, i.e., current RGB images $\mathbf{i}_{t}=[\mathbf{I}^{1}_{t},...,\mathbf{I}^{n}_{t}]$ at time step $t$, natural language instructions $\mathbf{\ell_{t}}$, and robot state information $\mathbf{s}_{t}$. In this way, the value-query head estimate the state-action value $Q_{\theta}(\mathbf{q}_{t},\mathbf{A}_{t}^{\tau_{n}})$ of the action chunk $\mathbf{A}_{t}^{\tau_{n}}$ conditioned on the input query token $\mathbf{q}_{t}$. We visualize the state-action value map of candidate actions in Fig. 7, and provide a detailed analysis in Sec. 6.2. This value-query head is trained on pre-collected robot demonstration dataset with ground-truth action $\mathbf{A}_{t}$ via offline RL levine2020offlinereinforcementlearningtutorial . We construct the training dataset $\mathcal{D}$ using the reward function following ptr , where the rewards of last 3 transitions in one robot episode is defined as $+1$, and the rest is 00. During training, we use the calibrated Q-learning algorithm nakamoto2023calql to optimize the value-query head.
@@
Value-Guided Thinking The System-2 value-guided thinking is implemented with Best-of-N selection strategy and the selected action chunk is conveyed to System 1 for cascaded action denoising. Specifically, conditioned on the same observation, the action denoising head generates $N$ candidate action chunks $\{\mathbf{A}_{t}^{\tau_{1}},\ \mathbf{A}_{t}^{\tau_{2}},\ \ldots,\ \mathbf{A}_{t}^{\tau_{N}}\}$ with different noise levels. Then, these candidates are passed to the value-query head to estimate their state-action values. Guided by the estimated state-action values, we select the action with the highest value as the optimal candidate $\mathbf{A}_{t}^{\tau^{*}}$ transferred to System 1, depicted as $\mathbf{A}_{t}^{\tau^{*}}=\mathop{\operatorname*{arg\,max}}\nolimits_{\mathbf{A}_{t}^{\tau_{i}}}Q(\mathbf{q_{t}},\mathbf{A}_{t}^{\tau_{i}})$, where $i=1\ldots N$.
@@ 3.2 Cascaded Dual-system Action Denoising
In order to achieve rapid, reactive robot control, System 1 module needs to be lightweight and fast in inference. In detail, System-1 consists of a DINOv2-small visual encoder and a lightweight transformer for cascaded action denoising. Given the selected candidate action chunk $\mathbf{A}_{t}^{\tau^{*}}$ from System 2, the System 1 module takes the observation $\tilde{\mathbf{o}}_{t+kh}$ (including current image $\mathbf{i}_{t}=\{\mathbf{I}_{t}^{1},..,\mathbf{I}_{t}^{n}\}$, robot state $\mathbf{s}_{t}$), and sub-action chunks $\tilde{\mathbf{A}}_{t}$ segmented from the selected candidate $\mathbf{A}_{t}^{\tau^{*}}$ as input, and produces refined robot actions by continuously denoising on sub-action chunks $\tilde{\mathbf{A}}_{t}$.
Specifically, at timestep $t$, the selected action chunk from System 2 is $\tilde{\mathbf{A}}_{t}=[\mathbf{a}_{t},\mathbf{a}_{t+1},...,\mathbf{a}_{t+H-1}]$, then $\tilde{\mathbf{A}}_{t}$ is segmented into $K:=H/h$ sub-action chunks $\{\tilde{\mathbf{A}}_{t},\tilde{\mathbf{A}}_{t+h},\ldots,\tilde{\mathbf{A}}_{t+(K-1)h}\}$ with a horizon of $h$. The System 1 sequentially performs cascaded denoising on these sub-action chunks with observation $\tilde{\mathbf{o}}_{t+kh}$. Note that System 1 is much faster than System 2, so that System 2 could finish cascaded denoising on all sub-action chunks before next action chunk $\tilde{\mathbf{A}}_{t}$ arriving. System 1 module is trained with the same flow matching loss using the action denoising head of System 2:
@@
Note that, during training and inference, the generated candidate action chunks from System 2 are not fully denoised, i.e., $\tilde{\mathbf{A}}_{t+kh}^{\tau^{*}}\neq\tilde{\mathbf{A}}_{t+kh}^{1}$, requiring the continuous denoising for accurate action prediction. Following the continuous denoising ho2021cascaded from image generation, System 1 refine the action by intergrating the learned vector field from $\omega=0$ to $\omega=1$. Instead of starting with random noise, for the $k$-th sub-action chunk the integration process of System 1 starting with the sub-action chunk $\tilde{\mathbf{A}}_{t+kh}^{\tau^{*}}$:
@@
where $\mathbf{v}_{\theta}\left(\tilde{\mathbf{A}}_{t+kh}^{\omega},\mathbf{o}_{t+kh}\right)$ is the vector field learned by System 1. With the same forward Euler method used in System-2 action denoising head, System 1 produces the final denoised action $\tilde{\mathbf{A}}_{t+kh}^{1}$ with 10 denoise steps (corresponding to $\delta=0.1$). After all K sub-action chunks have been processed by System 1, System 2 will generate a new selected action chunk ${\mathbf{A}}_{t+H}^{\tau^{*}}$, and System 1 will continue to refine segments from the new selected action chunk.
@@ 3.3 Training and Deployment Strategy
During the inference phase, the System 2 and System 1 modules are cooperating at asynchronous mechanism to boost up the overall control frequency. Specifically, at the initial timestep $t$, the action denoising head of System 2 generates $N=5$ multiple action chunks $\mathbf{A}_{0}^{\tau_{n}}\in\ \{\mathbf{A}_{0}^{\tau_{1}},\ \mathbf{A}_{0}^{\tau_{2}},\ \ldots,\ \mathbf{A}_{0}^{\tau_{N}}\}$ with horizon of $H=30$ as candidates at 4 Hz. Then the selected optimal action candidate $\mathbf{A}_{t}^{\tau^{*}}$ is stored in a shared queue. Then $\tilde{\mathbf{A}}_{t}^{\tau^{*}}$, a sub-action chunk with horizon $h=15$, is segmented from the first $h$ steps of $\mathbf{A}_{t}^{\tau^{*}}$ and passed into System 1. System 1 removes the remaining noise from $\tilde{\mathbf{A}}_{t}^{\tau^{*}}$, and produces the fully denoised action $\tilde{\mathbf{A}}_{t}$ at 6 Hz and execute all $h=15$ actions on real robot immediately, resulting in a overall 90 Hz robot action control frequency. After the robot executes all $K=H/h=2$ sub-action chunks in $\tilde{\mathbf{A}}_{t}$, System 1 repeatly get the newest selected action chunks from the shared queue for subsequent action denoising. Due the different working frequencies of System 2 and System 1, they asynchronously cooperate to achieve a balance between slow, human-like thinking and fast, reactive real robot control.
@@ 4.1 Multitask Evaluation on Simulation Benchmarks
Evaluation Results. Tab. 1 present the LIBERO (liu2023libero, ) experimental results. We observe that Hume can be effectively adapted to tasks in the LIBERO environments, as it obtains the highest average success rate of 98.6% and the first rank across all the policies. In particular, Hume achieves a remarkable 96.7% success rate (+11.5% over $\pi_{0}$, +6.1% over GR00T) on the LIBERO-Long task, which consists of long-horizon tasks, demonstrating the model’s strong long-term planning capabilities. Tab. 2 presents the SimplerEnv experimental results on WidowX and Google robot tasks. Hume also achieves state-of-the-art performance on WidowX multitasks, with an average success rate of 72.6%, representing a significant improvement compared to all current generalist manipulation policies (+32.5% over $\pi_{0}$, +39.6% over GR00T, +64.8% over OpenVLA). Similarly, Hume achieves an average success rate of 76.4% (+19.6% over $\pi_{0}$) on Google robot tasks. In summary, Hume demonstrates its versatility as a generalist robot control policy, achieving better performance across various tasks.
@@
Table 2: SimplerEnv evaluation across different policies on robot tasks. The zero-shot results denote performance of OXE dataset (o2024open, ) pre-trained models.
@@ 4.2 Real-World Embodiment Control
Real-world WidowX Evaluation. Fig. 3 presents the results of the real-world evaluation in WidowX robot platform. We compared some representative single-system VLA models and dual-system VLA models on multiple tasks. We observe that, in simple task scenarios (#1-2), most policies exhibit some generalizability, successfully completing tasks in unseen environments. However, in more complex tasks (#3-7), policies such as GR00T, $\pi_{0}$-FAST, and OpenVLA struggle with manipulation, frequently encountering grasp failures issues, such as inability to accurately grasp or place target objects. Even when the policies attempt to recover from failures, they often fall in error states that prevent successful execution. In contrast, Hume leverages value-guided thinking to effectively recover from failures, demonstrating superior performance on various complex tasks. That is, when Hume falls into a wrong state, it can evaluate multiple candidate actions and select the another correct trajectory forward. As a result, even if failures occur during initial attempts, Hume can adjust its trajectory and successfully complete the task on subsequent tries (please refer to supplementary video for more details), achieving strong robustness ability across various complex unseen tasks (91% average success rate), improves by +12% over $\pi_{0}$, and +33% over OpenVLA.
Real-world Franka and Agibot G-1 Evaluation. Fig. 4 presents the results of the real-world evaluation on Franka and Agibot G-1 robot platforms. The task design incorporates multiple real-world daily long-horizon tasks, deformable objects manipulation, tool-usage, and other challenging scenarios, with further validation conducted on both the Franka robot and the humanoid robot Agibot G-1. We observe that the long-term planning capability of System-2 thinking employed by Hume helps us better solve long-horizon tasks. For example, in Agibot’s long-horizon deformable objects manipulation task (#Fold Shorts), where the models need to make the robot fold two shorts, Hume achieves a success rate of 88%, improving by +15% over $\pi_{0}$. In the complex long-horizon task (#Pour Water), Hume achieves success rate of 82%, significantly improving by +20% over $\pi_{0}$, and +60% over GR00T. Additionally, Hume also achieves an average success rate of 87% across various tasks on the Franka robot, improving by +14.75% over $\pi_{0}$, and +37.25% over OpenVLA.
Figure 3: Real-world evaluation on WidowX Robot tasks. We evaluate Hume across 10 zero-shot tasks with varying backgrounds, poses, and motion distractors. Hume achieves the highest average success rate, surpassing $\pi_{0}$ and all other generalist manipulation policies in comparative evaluations.
@@ 4.3 Ablations on Design Decisions
Furthermore, we compared Hume with models directly using the candidate with the highest value output by System 2 (#1v.s.#4). Without System 1 to remove the remaining noise, models w/o System 1 cannot perform precise and dexterous movements, ultimately leading to an average decrease of -9.8% across SimplerEnv tasks, -8.8% across LIBERO tasks, and -63% in real-world robot tasks.
Effect of Value-Guided Thinking. According to the ablation results (#1v.s.#5), the proposed value-guided thinking enables the System 2 to select the most valuable candidate from multiple noisy candidates to pass to System 1, which effectively improves Hume’s performance in handling robot control tasks. Models w/o value-guided thinking randomly select 1 out of 5 candidate actions generated by System 2 to pass to System 1. Since there are candidates with different levels of noise in the candidate actions, this random selection strategy may select harmful candidates for System 1, ultimately resulting in a significant decrease in success rate, showing an average decline of -14.95% across multiple SimplerEnv tasks, -13.7% across LIBERO tasks. Notably, this performance degradation is even more pronounced in more complex real-world robot scenarios, with an astonishing average decrease of -78% across various tasks in real-world environments.

### v2

@@ 3.1 Value-Guided System-2 Thinking
State-Action Value Estimation The state-action values are estimated with the proposed value-query head built on the same VLM backbone via learning a latent conditioned Q function. The value-query head is composed of two critic networks estimating state-action values and one actor network for assisting the training of critic networks. Specifically, a special query token $\mathbf{q}_{t}$ is introduced and attached at the end of the VLM input sequence, which is a learnable token with the same embedding dimension as the language tokens. Then, for one action chunk $\mathbf{A}_{t}$ (either ground-truth action $\mathbf{A}_{t}^{\text{GT}}$ or candidate actions $\mathbf{A}_{t}^{\tau_{n}}$), it is combined with this special query token $\mathbf{q}_{t}$ to feed into the value-query head. Due to its last position of the input sequence, the query token $\mathbf{q}_{t}$ attends to all previous tokens and aggregates necessary information from the VLM inputs, i.e., current RGB images $\mathbf{i}_{t}=[\mathbf{I}^{1}_{t},...,\mathbf{I}^{n}_{t}]$ at time step $t$, natural language instructions $\mathbf{\ell_{t}}$. In this way, the value-query head estimate the state-action value $Q_{\theta}(\mathbf{q}_{t},\mathbf{A}_{t}^{\tau_{n}})$ of the action chunk $\mathbf{A}_{t}^{\tau_{n}}$ conditioned on the input query token $\mathbf{q}_{t}$. This value-query head is trained on pre-collected robot demonstration dataset with ground-truth action $\mathbf{A}_{t}$ via offline RL levine2020offlinereinforcementlearningtutorial . We construct the training dataset $\mathcal{D}$ using the reward function following ptr , where the rewards of last 3 transitions in one robot episode is defined as $+1$, and the rest is 00. During training, we use the calibrated Q-learning algorithm nakamoto2023calql to optimize the value-query head.
To verify our training pipeline can effectively optimize the value-query head, we project ground-truth actions $\mathbf{A}_{t}^{\text{GT}}$ and candidate actions $\mathbf{A}_{t}^{\tau_{n}}$ along with their corresponding state-action values $Q(\mathbf{q_{t}},\mathbf{A}_{t}^{\tau_{n}})$ into the same space, generating the value map in Fig. 2. The value map consists of regions representing different magnitudes of state-action values, and the ground-truth actions are all located in the high-value region, which proves the value-query head can estimate reasonable state-action value. The detailed analysis can be found in Sec. A.2
Figure 2: Value Map of Candidate Actions. The candidate actions $\mathbf{A}_{t}^{\tau_{n}}$ sampled from System 2 and ground-truth actions $\mathbf{A}_{t}^{\text{GT}}$ are projected into the same two-dimensional space through Principal Component Analysis (PCA). The intensity of colors indicates the magnitude of state-action values $Q(\mathbf{q_{t}},\mathbf{A}_{t}^{\tau_{n}})$ of candidate actions.
@@
Value-Guided Thinking The System-2 value-guided thinking is implemented with Best-of-N selection strategy and the selected action chunk is conveyed to System 1 for cascaded action denoising. Specifically, conditioned on the same observation, the action denoising head generates $N$ candidate action chunks $\{\mathbf{A}_{t}^{\tau_{1}},\ \mathbf{A}_{t}^{\tau_{2}},\ \ldots,\ \mathbf{A}_{t}^{\tau_{N}}\}$ with different noise levels. Then, these candidates are passed to the value-query head to estimate their state-action values. Guided by the estimated state-action values, we select the action with the highest value as the optimal candidate $\mathbf{A}_{t}^{\tau^{*}}$ transferred to System 1, depicted as $\mathbf{A}_{t}^{\tau^{*}}=\mathop{\operatorname*{arg\,max}}\nolimits_{\mathbf{A}_{t}^{\tau_{n}}}Q(\mathbf{q_{t}},\mathbf{A}_{t}^{\tau_{n}})$, where $n=1\ldots N$. In Sec. A.1, we further visualize this process in the simulated Push-T task.
@@ 3.2 Cascaded Dual-system Action Denoising
In order to achieve rapid, reactive robot control, System 1 module needs to be lightweight and fast in inference. In detail, System 1 consists of a DINOv2-small visual encoder and a lightweight transformer for cascaded action denoising. Given the selected candidate action chunk $\mathbf{A}_{t}^{\tau^{*}}$ from System 2, the System 1 module takes the observation $\tilde{\mathbf{o}}_{t+kh}$ (including current image $\mathbf{i}_{t}=\{\mathbf{I}_{t}^{1},..,\mathbf{I}_{t}^{n}\}$, robot state $\mathbf{s}_{t}$), and sub-action chunks $\tilde{\mathbf{A}}_{t}$ segmented from the selected candidate $\mathbf{A}_{t}^{\tau^{*}}$ as input, and produces refined robot actions by continuously denoising on sub-action chunks $\tilde{\mathbf{A}}_{t}$.
Specifically, at timestep $t$, the selected action chunk from System 2 is $\mathbf{A}_{t}^{\tau^{*}}=[\mathbf{a}_{t},\mathbf{a}_{t+1},...,\mathbf{a}_{t+H-1}]$, then $\mathbf{A}_{t}^{\tau^{*}}$ is segmented into $K:=H/h$ sub-action chunks $\{\tilde{\mathbf{A}}_{t},\tilde{\mathbf{A}}_{t+h},\ldots,\tilde{\mathbf{A}}_{t+(K-1)h}\}$ with a horizon of $h$. The System 1 sequentially performs cascaded denoising on these sub-action chunks with observation $\tilde{\mathbf{o}}_{t+kh}$. Note that System 1 is much faster than System 2, so that System 1 could finish cascaded denoising on all sub-action chunks before next action chunk $\mathbf{A}_{t}^{\tau^{*}}$ arriving. System 1 module is trained with the same flow matching loss using the action denoising head of System 2:
@@
Note that, the generated candidate action chunks from System 2 are not fully denoised, i.e., $\tilde{\mathbf{A}}_{t+kh}^{\tau^{*}}\neq\tilde{\mathbf{A}}_{t+kh}^{1}$, requiring the continuous denoising for accurate action prediction. Following the continuous denoising ho2021cascaded from image generation, System 1 refine the action by intergrating the learned vector field from $\omega=0$ to $\omega=1$. Instead of starting with random noise, the integration process of System 1 starting with the sub-action chunk $\tilde{\mathbf{A}}_{t+kh}^{\tau^{*}}$:
@@
where $\mathbf{v}_{\theta}\left(\tilde{\mathbf{A}}_{t+kh}^{\omega},\mathbf{o}_{t+kh}\right)$ is the vector field learned by System 1. Using the forward Euler method $\tilde{\mathbf{A}}_{t+kh}^{\omega+\sigma}=\tilde{\mathbf{A}}_{t+kh}^{\omega}+\delta\mathbf{v}_{\theta}\left(\tilde{\mathbf{A}}_{t+kh}^{\omega},\tilde{\mathbf{o}}_{t+kh}\right)$, System 1 produces the final denoised action $\tilde{\mathbf{A}}_{t+kh}^{1}$ with 10 denoise steps (corresponding to $\delta=0.1$). After all K sub-action chunks have been processed by System 1, System 2 will generate a new selected action chunk ${\mathbf{A}}_{t+H}^{\tau^{*}}$, and System 1 will continue to refine segments from the new selected action chunk.
@@ 3.3 Training and Deployment Strategy
During the inference phase, the System 2 and System 1 modules are cooperating at asynchronous mechanism to boost up the overall control frequency. Specifically, at the initial timestep $t$, the action denoising head of System 2 generates $N=5$ multiple action chunks $\mathbf{A}_{t}^{\tau_{n}}\in\ \{\mathbf{A}_{t}^{\tau_{1}},\ \mathbf{A}_{t}^{\tau_{2}},\ \ldots,\ \mathbf{A}_{t}^{\tau_{N}}\}$ with horizon of $H=30$ as candidates at 4 Hz. Then the selected optimal action candidate $\mathbf{A}_{t}^{\tau^{*}}$ is stored in a shared queue. Then $\tilde{\mathbf{A}}_{t}^{\tau^{*}}$, a sub-action chunk with horizon $h=15$, is segmented from the first $h$ steps of $\mathbf{A}_{t}^{\tau^{*}}$ and passed into System 1. System 1 removes the remaining noise from $\tilde{\mathbf{A}}_{t}^{\tau^{*}}$, and produces the fully denoised action $\tilde{\mathbf{A}}_{t}$ at 6 Hz and execute all $h=15$ actions on real robot immediately, resulting in a overall 90 Hz robot action control frequency. After the robot executes all $K=H/h=2$ sub-action chunks in $\tilde{\mathbf{A}}_{t}$, System 1 repeatly get the newest selected action chunks from the shared queue for subsequent action denoising. Due the different working frequencies of System 2 and System 1, they asynchronously cooperate to achieve a balance between slow, human-like thinking and fast, reactive real robot control.
@@ 4.1 Multitask Evaluation on Simulation Benchmarks
Evaluation Results. Tab. 1 present the LIBERO (liu2023libero, ) experimental results. We observe that Hume can be effectively adapted to tasks in the LIBERO environments, as it obtains the highest average success rate of 98.6% and the first rank across all the policies. In particular, Hume achieves a remarkable 96.7% success rate (+11.5% over $\pi_{0}$, +6.1% over GR00T) on the LIBERO-Long task, which consists of long-horizon tasks, demonstrating the model’s strong long-term planning capabilities. Tab. 2 presents the SimplerEnv experimental results on WidowX and Google robot tasks. Hume also achieves state-of-the-art performance on WidowX multitasks, with an average success rate of 72.6%, representing a significant improvement compared to all current generalist manipulation policies (+32.5% over $\pi_{0}$, +64.8% over OpenVLA). Similarly, Hume achieves an average success rate of 76.4% (+19.6% over $\pi_{0}$) on Google robot tasks. In summary, Hume demonstrates its versatility as a generalist robot control policy, achieving better performance across various tasks.
@@
Table 2: Evaluation results across different policies on SimplerEnv. We evaluate Hume on 4 tasks of WidowX and 3 tasks on Google-EDR in SimplerEnv
@@ 4.2 Real-World Embodiment Control
Real-world WidowX Evaluation. Fig. 4 presents the results of the real-world evaluation in WidowX robot platform. We compared some representative single-system VLA models and dual-system VLA models on multiple tasks. We observe that, in simple task scenarios (#1-2), most policies exhibit some generalizability, successfully completing tasks in unseen environments. However, in more complex tasks (#3-7), policies such as GR00T, $\pi_{0}$-FAST, and OpenVLA struggle with manipulation, frequently encountering grasp failures issues, such as inability to accurately grasp or place target objects. Even when policies attempt to recover from failures, they often trap in error states that prevent successful execution. In contrast, Hume leverages value-guided thinking to effectively recover from failures, demonstrating superior performance on various complex tasks. That is, when Hume falls into a wrong state, it can evaluate multiple candidate actions and select another correct trajectory forward. As a result, even if failures occur during initial attempts, Hume can adjust its trajectory and successfully complete the task on subsequent attempts (please refer https://hume-vla.github.io for more details), achieving strong robustness ability across various complex unseen tasks (91% average success rate), improves by + 12% over $\pi_{0}$ and + 33% over OpenVLA.
Figure 4: Real-world evaluation on WidowX Robot tasks. We evaluate Hume across 10 tasks with varying backgrounds, poses, and motion distractors. Hume achieves the highest average success rate, surpassing $\pi_{0}$ and all other generalist manipulation policies in comparative evaluations.
Real-world Franka and Agibot G-1 Evaluation. Fig. 5 presents the results of the real-world evaluation on Franka and Agibot G-1 robot platforms. The task design incorporates multiple real-world long-horizon tasks, deformable objects manipulation, tool-usage, and other challenging scenarios, with further validation conducted on both the Franka robot and the humanoid robot Agibot G-1. We observe that the long-term planning capability of System-2 thinking employed by Hume helps us better solve long-horizon tasks. For example, in Agibot’s long-horizon deformable objects manipulation task (#Fold Shorts), where the models need to make the robot fold different shorts, Hume achieves a success rate of 88%, improving by +15% over $\pi_{0}$. In the complex long-horizon task (#Pour Water), Hume achieves success rate of 82%, significantly improving by +20% over $\pi_{0}$, and +60% over GR00T. Additionally, Hume also achieves an average success rate of 87% across various tasks on the Franka robot, improving by +14.75% over $\pi_{0}$, and +37.25% over OpenVLA.
@@ 4.3 Ablations on Design Decisions
Furthermore, we compared Hume with the model directly using the optimal candidate generated by System 2 (#1v.s.#4). Without System 1 to remove the remaining noise, the model w/o System 1 cannot perform precise and dexterous movements, ultimately leading to an average decrease of -9.8% across SimplerEnv tasks, -8.8% across LIBERO tasks, and -63% in real-world robot tasks.
Effect of Value-Guided Thinking. According to the ablation results (#1v.s.#5), the proposed value-guided thinking enables the System 2 to select the most valuable candidate from multiple noisy candidates to pass to System 1, which effectively improves Hume’s performance in handling complex robot control tasks. Models w/o value-guided thinking randomly select 1 out of 5 candidate actions generated by System 2 to pass to System 1. Since there are candidates with different levels of noise in the candidate actions, this random selection strategy may select harmful candidates for System 1, ultimately resulting in a significant decrease in success rate, showing an average decline of -14.95% across multiple SimplerEnv tasks, -13.7% across LIBERO tasks. Notably, this performance degradation is even more pronounced in more complex real-world robot scenarios, with an astonishing average decrease of -78% across various tasks in real-world environments.